OpenStudy (anonymous):

6. If the arc length shown is 18.75 cm and the radius of the circle is 5 cm, find theta to the nearest hundredth of a degree

OpenStudy (anonymous):

18.75 = 5 x theta => theta = 18..75 / 5

OpenStudy (wolf1728):

arc length / radius = central angle 18.75 / 5 = central angle (radians) 18.75 / 5 = 3.75 radians which equals 214.8591731741 degrees or 214.86 degrees (rounded to nearest hundredth)

OpenStudy (anonymous):

thanks guys ive got it figured out but thanks for you're help :)

OpenStudy (wolf1728):

u r welcome